OpenStudy (anonymous):

traffic signs are regulated by the Manual on Uniform traffic control devices ( MUTCD) The perimeter of some traffic sign must be 128 inches, also, its lenght is 10inch longer than its width. find the dimensions of this sign

OpenStudy (stormfire1):

For perimeter: P = 2L + 2W L = W + 10 128 = 2L + 2W 64 = L + W 64 = (W+10) + W 54 = 2W W = 27 So the width will be 27 inches and the length will be 27 + 10 = 37 inches.